OpenStudy (anonymous):

can someone please help me solve this equation? I am new to this but am just looking for help in understanding. f/7 -3 =1

OpenStudy (anonymous):

If you equation is: \(\dfrac{f}{7}-3=1\) Use inverse operations to isolate the variable. What you do to one side you must do to the other to keep the equation balanced. Add 3 to both sides. \(\dfrac{f}{7}-3+3=1+3\implies \dfrac{f}{7}=4\) Multiply both sides by 7: \(\dfrac{f}{\cancel7}\times \cancel7 = 4 \times 7 \implies f=28\)